OpenStudy (anonymous):

Chain Rule Differentiation! \(\ \Huge F(x)=\sqrt[4]{1+2x+x^3} . \)

OpenStudy (anonymous):

\(\ \Huge F(x)=\sqrt[4]{1+2x+x^3} . \) \[\ \large \text{The answer I got was:} \frac{1}{4}(x^3+2x+1) \times (3x^2+2) .\] \[\ \large \text{Is this correct?} \]

zepdrix (zepdrix):

So when you applied the power rule, you changed it to a 1/4 power, but then you didn't subtract 1 from the power, you actually increased it by 3/4. The new power should be 1/4 - 1, simply that :D

OpenStudy (anonymous):

\(\ \text{Oh that's a typo on my part, I have} -\frac{3}{4} \text{as my final answer.}\)

zepdrix (zepdrix):

Oh cool c: yah looks good!

OpenStudy (anonymous):

\(\ \large \text{Thanks @zepdrix for checking my work! :)} \)

zepdrix (zepdrix):

How do you type in that fancy text? :C hmm

zepdrix (zepdrix):

\text{test}\

zepdrix (zepdrix):

grr :d

OpenStudy (anonymous):

\(\ \Huge \text{with } LATEX, \text{@zepdrix}... \)

zepdrix (zepdrix):

\( \text{Interesting! :O I gosta learn that!} \)

OpenStudy (anonymous):

\(\ \Huge \text{:D} \)